I can take this so far then I get lost can you help me figure it out? Soybean meal is 16% protein and cornmeal is 8% protein how much of each would it take to make a 320 pound mixture that is 10% protein
Answer by stanbon(75887)

Soybean meal is 16% protein and cornmeal is 8% protein how much of each would it take to make a 320 pound mixture that is 10% protein
---------------------------
Equation:
protein + protein = protein
0.16x + 0.08(320-x) = 0.10*320
---
Multiply thru by 100 to get:
16x + 8*320 - 8x = 10*320
8x = 2*320
x = 80 lbs (amt. of 16% meal needed in the mixture)
---
320-x = 240 (amt. of 8% meal needed in the mix)
=====================================================
Cheers,
Stan H.